OpenStudy (anonymous):

negative 1 over 2 x + 1 = −x + 8

OpenStudy (anonymous):

take value of x at one side and other on other side of equality

OpenStudy (anonymous):

\[{-1 \over 2x+1} =-x+8\] -1 = (-x+8)(2x+1) -1= -2x^2 -x +16x+8

OpenStudy (anonymous):

start with cross multiple

OpenStudy (anonymous):

Then 0= -2x^2 +15x +9 0= 2x^2-15x -9

OpenStudy (anonymous):

Then use quadratic equation?

OpenStudy (anonymous):

oh